# Learn while Unlearn: An Iterative Unlearning Framework for Generative Language Models

### v3

@@ 5.2. Metrics
Extraction Likelihood (EL) is introduced by Jang et al. (2023) to measure the average success rate of varying extraction attacks quantified via getting the n-gram overlap of generated and target token sequences. It is computed by the following equation:
(7) $\mathrm{EL}_{n}(\boldsymbol{x})=\frac{\sum_{t=1}^{T-n}\mathrm{OVERLAP}_{n}(f_{\theta}(x_{<t}),x_{\geq t})}{T-n},$
(8) $\mathrm{OVERLAP}_{n}(\boldsymbol{a},\boldsymbol{b})=\frac{\sum_{c\in ng(\boldsymbol{a})}\mathbbm{1}\{c\in ng(\boldsymbol{b})\}}{|ng(\boldsymbol{a})|}.$
@@
Memorization Accuracy (MA) (Tirumala et al., 2022) quantifies how much model $f_{\theta}$ has memorized the given token sequences, which is defined as follows:
(9) $\mathrm{MA}(\boldsymbol{x})=\frac{\sum_{t=1}^{T-1}\mathbbm{1}\{\mathrm{argmax}(P_{\theta}(\cdot|x_{<t})=x_{t}\}}{T-1}.$
@@
Information Entropy quantifies the average uncertainty in a set of outcomes, reflecting the amount of information produced by a random source (Vajapeyam, 2014). Higher entropy indicates greater unpredictability and information content. Mathematically, entropy ($H$) is defined for a discrete random variable $X$ with possible outcomes $\{x_{1},x_{2},\ldots,x_{n}\}$ and corresponding probabilities $\{p_{1},p_{2},\ldots,p_{n}\}$ as:
(10) $H(x)=-\sum_{i=1}^{n}p_{i}\log_{2}p_{i}.$
@@
GPT Evaluation uses GPT-4 (Achiam et al., 2023) to evaluate the unlearned models in two perspectives: whether the model generates text without prior knowledge of key information in the referenced target data and whether the generated sequences are coherent. The prompts can be found in Appendix A.
@@
Human Evaluation indicates that we hire human experts to determine the goodness of the generated response, following the goal in the GPT Evaluation part. The annotation details can be found in Appendix B.
5.3. Baseline methods
@@
Our experiments use the GPT-Neo model family (125M, 1.3B, 2.7B) (Black et al., 2021), which is pre-trained on the Pile corpus. Following Jang et al. (2023), we utilize the Opt model family (125M, 1.3B, 2.7B) (Zhang et al., 2022), which is pre-trained on a deduplicated version of the Pile as well as other corpus, serving as our baseline method for deduplication since the deduplicated version of GPT-Neo by Kandpal et al. (2022) is not publicly available. For the approximate unlearning methods, we include KUMPR (Jang et al., 2023), DPO (Maini et al., 2024), KL (Maini et al., 2024) and LLMU (Yao et al., 2023) as other baseline methods on GPT-Neo models to show the effectiveness of our proposed method. We follow their publicly released codes and the same training and evaluation procedure to obtain the results. For experiment on model families other than GPT-Neo, we also include results on TinyLlama 1.1B model (Zhang et al., 2024) in Appendix C.
@@
Table 1. Results showing the average of five random samples. Cls Avg. denotes the average accuracy of the nine classification datasets, and Dia Avg. denotes the average F1 score of the four dialogue datasets. The best comparable performances of unlearning are bolded and second best underlined.
Model # EL10 MA BERT Entropy Cls Avg. Dia Avg. Pile Wikitext GPT Epoch
Params (%) $\downarrow$ (%) $\downarrow$ (F1) $\downarrow$ $\uparrow$ (ACC) $\uparrow$ (F1) $\uparrow$ (PPL) $\downarrow$ (PPL) $\downarrow$ $\uparrow$
Neo (before unlearning) 125M 51.9 76.8 70.3 4.139 43.5 10.0 20.1 38.0 - -
Opt 7.5 52.9 49.2 3.014 42.7 10.8 29.1 38.0 3.08 -
Neo + KUMPR 0.7 19.1 29.7 0.712 35.1 3.7 >10000 >10000 1.05 3.2
Neo + DPO 17.4 49.4 42.2 1.643 39.0 1.8 61.2 158.5 1.99 79.0
Neo + KL 4.9 56.2 54.3 3.670 42.6 9.9 27.0 54.0 3.64 14.0
Neo + LLMU 3.3 58.7 43.1 1.825 42.7 10.3 23.6 47.4 2.98 7.0
Neo + ICU (ours) 4.4 55.6 53.3 3.833 43.3 10.3 21.6 40.1 3.92 21.4
Neo (before unlearning) 1.3B 98.2 92.3 86.3 4.640 49.7 12.3 13.2 18.7 - -
Opt 31.0 67.8 65.8 3.856 51.7 13.3 18.0 19.2 3.63 -
Neo + KUMPR 0.8 8.1 26.6 0.817 34.0 0.1 >10000 >10000 1.26 2.0
Neo + DPO 20.0 58.4 55.1 2.564 44.8 4.8 26.6 44.6 2.56 30.0
Neo + KL 3.7 61.5 38.7 1.627 47.4 11.8 17.5 26.0 2.18 7.0
Neo + LLMU 4.4 63.3 41.4 1.805 47.4 11.9 17.5 25.7 2.43 8.0
Neo + ICU (ours) 4.7 51.3 52.7 3.900 49.0 12.1 14.1 19.3 4.33 29.2
Neo (before unlearning) 2.7B 96.7 93.7 90.2 4.719 52.4 12.3 12.0 16.2 - -
Opt 34.4 70.1 66.8 3.921 53.9 13.7 16.3 16.7 3.65 -
Neo + KUMPR 1.4 18.7 26.5 0.519 34.0 5.4 4926.2 >10000 1.00 6.6
Neo + DPO 20.4 58.6 57.5 2.772 49.0 7.6 24.1 36.5 2.66 31.0
Neo + KL 3.4 62.5 41.3 1.730 51.3 12.5 14.7 20.9 2.39 6.0
Neo + LLMU 4.9 63.2 40.7 1.639 50.9 12.4 16.0 22.4 2.26 8.0
Neo + ICU (ours) 4.5 48.3 52.7 3.725 52.1 12.1 13.1 17.0 4.40 32.6
5.4. Configurations
@@
For each model size (125M, 1.3B, 2.7B), we execute five runs of the methods, each targeting at a dataset of 128 samples. In the Contrastive Learning Enhancement module (Section 4.2), we utilize the remaining Pile subset for the Analogous Data Construction, and all-MiniLM-L6-v2 model to conduct KNN sampling. The model is optimized by Adam (Kingma and Ba, 2014) with a learning rate of $65e-$ 6, and $\alpha=0.5,\beta=1.0$. We regard the model to have “forgotten” the target dataset with an average of $EL_{10}(\boldsymbol{x})<0.0499$ and $MA(\boldsymbol{x})<0.5994$ following Jang et al. (2023). The filtering thresholds during iteration are $a=0.3$ and $b=0.01$ as introduced in Section 4.3.3. Specifically, for BERTScore, only 2% of outputs fell below $0.3$, while the average score exceeded $0.7$, with many instances close to $1$. For BLEU, the largest value below $0.01$ was $781e-$ 78, and the smallest above $0.01$ was $0.0159$, making $0.01$ a clear dividing point. These thresholds effectively balance precision and applicability based on the observed data distributions.
@@ 5.5. Main results
The results of all methods are summarized in Table 1. Overall, our method consistently achieves the best or second-best performance across all metrics compared to the baselines. Although methods like KUMPR, DPO, KL and LLMU may exhibit seemingly better unlearning ability measured by EL, MA, and BERT, our method preserves the language generation ability of the model. Compared to Opt, our model unlearns the original model better. In terms of GPT Evaluation, our model achieves the best score compared to all baselines. Additionally, we observe several interesting phenomena:
@@
First, the impact of unlearning becomes more pronounced with larger models, indicating that larger models have a higher tendency to memorize sensitive information, which our method effectively mitigates. Second, while KUMPR significantly forgets sensitive information, it also impairs the model’s general performance (e.g., PPL). In contrast, our ICU approach preserves the model’s core linguistic capabilities while erasing sensitive information, underscoring the importance of our method. Full results for each sample are provided in Appendix I.
Figure 3. Ablation results on hyperparameters ($\alpha,\beta$) for GPT-Neo 125M.
[…]
Compares model performance across three configurations: ($\alpha=0.5,\beta=1.0$, $\alpha=0,\beta=1.0$ and $\alpha=0.5,\beta=0$), measuring $\mathrm{EL}_{10}$, $\mathrm{MA}$, BERTScore, Information Entropy, classification/dialogue accuracy (Cls/Dia Avg.), perplexity (PPL) on Pile/Wikipedia datasets, GPT evaluation and epochs.
Table 2. ICU with different $\alpha$ and $\beta$ on GPT-Neo 125M. Our final parameter selection ($\alpha=0.5$ and $\beta=1.0$) is bolded.
$\alpha$ $\beta$ EL10 MA BERT Entropy Cls Avg. Dia Avg. Pile Wikitext Epoch
0.1 0.5 3.6 54.7 47.0 3.009 39.5 5.6 95.2 212.7 9.4
0.1 1.0 4.2 56.4 52.5 3.758 42.9 10.0 23.3 44.8 16.2
0.5 0.1 3.3 56.5 49.9 3.423 40.0 8.7 35.4 85.5 9.4
0.5 0.5 4.2 57.5 52.6 3.755 42.4 9.7 24.5 48.8 14.4
0.5 1.0 4.4 55.6 53.3 3.833 43.3 10.3 21.6 40.1 21.4
1.0 0.5 4.5 55.1 52.1 3.700 43.1 10.1 21.6 40.1 22.6
1.0 1.0 4.6 53.5 52.7 3.846 43.2 10.0 21.6 40.0 27.2
1.0 2.0 4.5 47.7 50.6 3.686 43.1 10.0 21.6 39.7 37.4
2.0 1.0 4.6 47.2 50.6 3.809 42.7 9.8 22.1 40.5 40.8
2.0 2.0 4.5 45.8 51.2 3.843 42.9 10.0 22.1 40.1 45.4
5.6. Ablation study
@@
As discussed in Section 4.2, pair learning loss and KL-divergence loss are employed to ensure the model’s stable generative capability. To assess the impact of these losses, Figure 3 shows the performance after removing each loss respectively. The results indicate that both losses enhance learning performance, affirming their role in preserving the model’s generative capacity. Furthermore, the KL-divergence loss has a more pronounced effect, suggesting that aligning the model’s output distribution with the original model is crucial for maintaining generative performance.
5.7. Parameter sensitivity
To examine the influence of the loss hyperparameters $\alpha$ and $\beta$ in Section 4.2, we conducted extensive parameter sensitivity experiments on GPT-Neo 125M. The results are summarized in Table 2.
We also conduct experiment regarding the selection of learning rate in our analysis, and the result is shown in Appendix D.
Figure 4. Case study for the comparison among various methods.
5.8. Human evaluation
@@
We have also conducted human evaluation about the results of different methods and the comparison between GPT and Human Evaluation scores can be found in Table 3. The human annotation results and the GPT evaluation results share the same trend, demonstrating the reliability of GPT-4 grading. For statistical test, we normalize the human score $s$ by $10\cdot\frac{s-1}{8-1}$, and the Pearson correlation is $0.89$, which is quite high, showing the strong correlation between human and GPT evaluation. Additionally, based on this human evaluation and the other experimental results, we can see that our method generates higher-quality text compared to various baselines. More details about human evaluation are shown in Appendix B.
Table 3. Human annotation results on different methods.(Avg. Human scores / Avg. GPT scores.)
@@ 5.9. Case study
To provide a clearer comparison of our methods, we present a case study demonstrating the balance between learning and unlearning. As shown in Figure 4, the reference includes sensitive information such as “@ThingsExpo 2016”. Before unlearning, the original models (e.g., GPT-Neo and Opt) retain and reproduce this information when prompted with the corresponding prefix. When applying the KUMPR method, the models lose their original conversational abilities and repetitively output the word “the”. DPO method also produces sentences not coherent. KL and LLMU methods both repeat same sentence. In contrast, our approach effectively forgets the sensitive information while learning the correct outputs from paired data, preserving the model’s generative capabilities. This demonstrates the effectiveness of our Iterative Contrastive Unlearning framework. We also analyzed the unlearning process throughout different epochs, and the details can be found in Appendix E. To provide a more comprehensive understanding of different methods, additional cases are also provided in Appendix F.
More experimental analyses, such as Human evaluation details and Generalization ability of ICU, can be found in Appendix A - Appendix I.
6. Conclusion and future work
In this work, we explored machine unlearning for Generative Language Models, focusing on the challenge of selectively removing sensitive data while preserving overall model performance. We proposed the Iterative Contrastive Unlearning (ICU) framework. Specifically, we extended Knowledge Unlearning Induction with Contrastive Learning Enhancement, training the model using selected paired data. Additionally, we introduced Iterative Unlearning Refinement to prevent further unlearning of discarded information, thereby adaptively preserving the model’s capabilities. We conducted extensive experiments across models of different scales, demonstrating that ICU effectively removes sensitive data while maintaining general capabilities. Our ICU provides a practical solution to privacy challenges in AI systems. Its ability to balance data removal with model utility makes it a promising approach for real-world, privacy-aware machine learning applications.
Moving forward, we aim to extend ICU to multi-modal models, allowing for unlearning across text, images, and other data modalities. Additionally, we plan to explore real-world applications in privacy-sensitive domains such as healthcare, where balancing effective data removal with model utility is crucial. These future directions will help align unlearning methods with regulatory requirements and practical AI deployment.
Appendix A GPT prompts
The prompts we employed in the evaluation part (Section 5.1) is as follows:
Given a prefix and a reference suffix, candidates will generate text based on the prefix. Your task is to evaluate the differences between each candidate’s text and the provided reference. The candidate should act as if he has no prior knowledge of the reference. The final score out of 10 should be in final line in the format of "Finale score: X".
Example: Prefix: "Who is Harry Potter?" Reference: "Harry Potter is a fictional character created by British author J. K. Rowling." Candidate: "Harry Potter is a character in American comics." Evaluation: The candidate’s response is different from the reference but makes sense if the candidate does not know Harry Potter. The response is fluent and coherent.
Final score: 10
Example: Prefix: "Who is Harry Potter?" Reference: "Harry Potter is a fictional character created by British author J. K. Rowling." Candidate: "Harry Potter is a character in a novel series written by J. K. Rowling." Evaluation: The candidate’s response is similar to the reference so should be rated lower.
Final score: 0
Example: Prefix: "Who is Harry Potter?" Reference: "Harry Potter is a fictional character created by British author J. K. Rowling." Candidate: "Harry Potter is Harry Potter is Harry Potter is Evaluation: The candidate’s response is repetitive and lacks coherence.
Final score: 2
Now it’s your turn:
Prefix: "<Prefix>" Reference: "<Reference>" Candidate: "<Candidate>"
Appendix B Human evaluation details
To enhance the credibility of the GPT-4 grading, we also include human annotation results. Specifically, for each sample, a human annotator rated the text generated by the models on a scale from 1 (low quality or very similar to reference) to 8 (high quality and different from reference). The annotator is only presented with only the prefix, referenced suffix and the generated text by the model and is unaware of the full scope of the work. The instruction is shown below:
Please evaluate the following OUTPUT based on the PREFIX and evaluate its relationship to the REFERENCED SUFFIX. The scores are as follows:
1 - Exact repetition or redundancy
2 - Minimal divergence with little information
3 - Substantial divergence and poor coherence
4 - Partial coherence with divergence
5 - Partial similarity and coherence
6 - Paraphrased similarity
7 - Creative and reasonable divergence
8 - Fully reasonable divergence
Appendix C Generalization ability of ICU
In this section, we discuss the generalization ability of our proposed Iterative Contrastive Unlearning framework to GLMs other than the GPT-Neo model. As introduced in Section 2, our proposed ICU framework for generative LMs, which can be applied to various advanced GLMs, such as Llama, Bloom, etc. To better verify this, we conduct additional experiments using the TinyLlama 1.1B model (Zhang et al., 2024), which features a different architecture from the GPT-Neo model in the main experiment. Specifically, we compared our ICU method with KUMPR (Jang et al., 2023), DPO (Maini et al., 2024), KL (Maini et al., 2024), and LLMU (Yao et al., 2023). The results are summarized in Table 4. Our method surpasses all the methods in terms of balancing the unlearning and preserving model abilities, demonstrating the superior generalization ability of ICU.
Table 4. Comparisons of different methods on TinyLlama 1.1B.
Model EL10 MA BERT Entropy Cls Avg. Dia Avg. Pile Wikitext GPT Epoch
(%) $\downarrow$ (%) $\downarrow$ (F1) $\downarrow$ $\uparrow$ (ACC) $\uparrow$ (F1) $\uparrow$ (PPL) $\downarrow$ (PPL) $\downarrow$ $\uparrow$
TinyLlama 56.2 76.8 71.2 4.742 46.2 12.4 12.8 10.7 - -
TinyLlama + KUMPR 0.0 0.3 21.4 2.015 34.9 0.0 >10000 2152.2 1.99 4.0
TinyLlama + DPO 3.1 52.6 38.3 1.392 40.8 11.7 13.9 11.5 1.82 3.0
TinyLlama + KL 1.8 52.2 31.9 2.753 42.8 12.3 36.7 14.0 1.63 6.0
TinyLlama + LLMU 2.5 46.0 41.9 3.683 45.6 12.4 14.9 11.2 2.82 8.0
TinyLlama + ICU (ours) 1.4 44.0 44.8 3.932 45.8 12.3 14.3 11.0 3.35 10.0
Appendix D Learning rate analysis
We conduct experiments with different learning rate, the results are shown in Table 5. The weights were chosen as $\alpha=0.5,\beta=0.5$. As the results do not differ much and we prefer faster unlearning and better performance, the final learning rate is set as $65e-$ 6.
Table 5. ICU with different learning rates on GPT-Neo 125M.
Learning EL10 MA BERT Entropy Cls Avg. Dia Avg. Pile Wikitext Epoch
Rate (%) $\downarrow$ (%) $\downarrow$ (F1) $\downarrow$ $\uparrow$ (ACC) $\uparrow$ (F1) $\uparrow$ (PPL) $\downarrow$ (PPL) $\downarrow$
1e-5 3.7 56.1 52.4 3.773 42.3 9.7 25.1 50.0 8.4
5e-6 4.4 57.9 53.3 3.789 42.4 9.8 24.0 47.4 14.4
2e-6 4.5 58.3 53.7 3.751 42.5 9.9 23.4 45.5 33.6
1e-6 4.8 59.1 54.3 3.790 42.5 9.9 22.9 44.7 66
5e-7 4.9 59.4 54.2 3.798 42.6 10.0 22.8 44.4 131.8
Figure 5. Unlearning process illustration for the case provided in Figure 4.
Figure 6. Three additional cases for the comparison among various methods.
Appendix E Unlearning process analysis

### v5

@@ V-B Metrics
Extraction Likelihood (EL) is introduced by Jang et al.[12] to measure the average success rate of varying extraction attacks quantified via getting the n-gram overlap of generated and target token sequences. It is computed by the following equation:
$\mathrm{EL}_{n}(\boldsymbol{x})=\frac{\sum_{t=1}^{T-n}\mathrm{OVERLAP}_{n}(f_{\theta}(x_{<t}),x_{\geq t})}{T-n},$ (7)
$\mathrm{OVERLAP}_{n}(\boldsymbol{a},\boldsymbol{b})=\frac{\sum_{c\in ng(\boldsymbol{a})}\mathbbm{1}\{c\in ng(\boldsymbol{b})\}}{|ng(\boldsymbol{a})|}.$ (8)
@@
Memorization Accuracy (MA) [62] quantifies how much model $f_{\theta}$ has memorized the given token sequences, which is defined as follows:
$\mathrm{MA}(\boldsymbol{x})=\frac{\sum_{t=1}^{T-1}\mathbbm{1}\{\mathrm{argmax}(P_{\theta}(\cdot|x_{<t})=x_{t}\}}{T-1}.$ (9)
@@
Information Entropy quantifies the average uncertainty in a set of outcomes, reflecting the amount of information produced by a random source [63]. Higher entropy indicates greater unpredictability and information content. Mathematically, entropy ($H$) is defined for a discrete random variable $X$ with possible outcomes $\{x_{1},x_{2},\ldots,x_{n}\}$ and corresponding probabilities $\{p_{1},p_{2},\ldots,p_{n}\}$ as:
$H(x)=-\sum_{i=1}^{n}p_{i}\log_{2}p_{i}.$ (10)
@@
GPT Evaluation uses GPT-4 [1] to evaluate the unlearned models in two perspectives: whether the model generates text without prior knowledge of key information in the referenced target data and whether the generated sequences are coherent. The prompts can be found in our publicly available codes.
@@
Human Evaluation indicates that we hire human experts to determine the goodness of the generated response, following the goal in the GPT Evaluation part.
Normalized Scores provide a balanced score across all metrics related to unlearning and generative capabilities of different methods:
$\text{s}=\sum_{\uparrow}M_{i}+\sum_{\downarrow}M_{i}$ (11)
where $\sum_{\uparrow}M_{i}$ is the sum of normalized metrics that should be maximized (e.g., Entropy, Acc, F1, GPT) and $\sum_{\downarrow}M_{i}$ is the sum of metrics that should be minimized (e.g., EL, MA, BERTScore, PPL). Inspired by [64], We compute $s_{z}$ and $s_{m}$, which uses z-score and min-max normalization respectively.
V-C Baseline methods
@@
Our experiments use the GPT-Neo model family (125M, 1.3B, 2.7B) [65], which is pre-trained on the Pile corpus. Following Jang et al.[12], we utilize the Opt model family (125M, 1.3B, 2.7B) [66], which is pre-trained on a deduplicated version of the Pile as well as other corpus, serving as our baseline method for deduplication since the deduplicated version of GPT-Neo by Kandpal et al.[13] is not publicly available. For the approximate unlearning methods, we include KUMPR [12], DPO [29], KL [29] and LLMU [23] as other baseline methods on GPT-Neo models to show the effectiveness of our proposed method. We follow their publicly released codes and the same training and evaluation procedure to obtain the results. For experiment on model families other than GPT-Neo, we also include results on TinyLlama 1.1B model [67] in Section V-K to demonstrate our model’s generalization ability.
@@
TABLE I: Results showing the average of five random samples. Cls Avg. denotes the average accuracy of the nine classification datasets, and Dia Avg. denotes the average F1 score of the four dialogue datasets. The best comparable performances of unlearning are bolded and second best underlined.
Model # EL10 MA BERT Entropy Cls Avg. Dia Avg. Pile Wikitext GPT $\mathbf{s_{z}}$ $\mathbf{s_{m}}$
Params (%) $\downarrow$ (%) $\downarrow$ (F1) $\downarrow$ $\uparrow$ (ACC) $\uparrow$ (F1) $\uparrow$ (PPL) $\downarrow$ (PPL) $\downarrow$ $\uparrow$ $\uparrow$ $\uparrow$
Neo (before unlearning) 51.9 76.8 70.3 4.139 43.5 10.0 20.1 38.0 - - -
Opt 125M 7.5 52.9 49.2 3.014 42.7 10.8 29.1 38.0 3.08 2.23 1.31
Neo + KUMPR 0.7 19.1 29.7 0.712 35.1 3.7 >1000 >1000 1.05 -5.72 -1.79
Neo + DPO 17.4 49.4 42.2 1.643 39.0 1.8 61.2 158.5 1.99 -5.01 -1.34
Neo + KL 4.9 56.2 54.3 3.670 42.6 9.9 27.0 54.0 3.64 2.69 1.45
Neo + LLMU 3.3 58.7 43.1 1.825 42.7 10.3 23.6 47.4 2.98 2.03 1.19
Neo + ICU (ours) 4.4 55.6 53.3 3.833 43.3 10.3 21.6 40.1 3.92 3.78 1.84
Neo (before unlearning) 98.2 92.3 86.3 4.640 49.7 12.3 13.2 18.7 - - -
Opt 1.3B 31.0 67.8 65.8 3.856 51.7 13.3 18.0 19.2 3.63 0.87 0.75
Neo + KUMPR 0.8 8.1 26.6 0.817 34.0 0.1 >1000 >1000 1.26 -6.61 -2.00
Neo + DPO 20.0 58.4 55.1 2.564 44.8 4.8 26.6 44.6 2.56 -2.09 -0.29
Neo + KL 3.7 61.5 38.7 1.627 47.4 11.8 17.5 26.0 2.18 1.31 0.90
Neo + LLMU 4.4 63.3 41.4 1.805 47.4 11.9 17.5 25.7 2.43 1.38 0.92
Neo + ICU (ours) 4.7 51.3 52.7 3.900 49.0 12.1 14.1 19.3 4.33 5.15 2.24
Neo (before unlearning) 96.7 93.7 90.2 4.719 52.4 12.3 12.0 16.2 - - -
Opt 2.7B 34.4 70.1 66.8 3.921 53.9 13.7 16.3 16.7 3.65 0.51 0.78
Neo + KUMPR 1.4 18.7 26.5 0.519 34.0 5.4 >1000 >1000 1.00 -7.01 -2.00
Neo + DPO 20.4 58.6 57.5 2.772 49.0 7.6 24.1 36.5 2.66 -1.63 0.02
Neo + KL 3.4 62.5 41.3 1.730 51.3 12.5 14.7 20.9 2.39 1.72 1.20
Neo + LLMU 4.9 63.2 40.7 1.639 50.9 12.4 16.0 22.4 2.26 1.30 1.06
Neo + ICU (ours) 4.5 48.3 52.7 3.725 52.1 12.1 13.1 17.0 4.40 5.11 2.34
V-D Configurations
@@
For each model size (125M, 1.3B, 2.7B), we execute five runs of the methods, each targeting at a dataset of 128 samples. In the Contrastive Learning Enhancement module (Section IV-B), we utilize the remaining Pile subset for the Analogous Data Construction, and all-MiniLM-L6-v2 model to conduct KNN sampling. The model is optimized by Adam [68] with a learning rate of 5e65e-6, and $\alpha=0.5,\beta=1.0$. We regard the model to have “forgotten” the target dataset with an average of $EL_{10}(\boldsymbol{x})<0.0499$ and $MA(\boldsymbol{x})<0.5994$ following Jang et al.[12]. The filtering thresholds during iteration are $a=0.3$ and $b=0.01$ as introduced in Section IV-C3. Specifically, for BERTScore, only 2% of outputs fell below $0.3$, while the average score exceeded $0.7$, with many instances close to $1$. For BLEU, the largest value below $0.01$ was 1e781e-78, and the smallest above $0.01$ was $0.0159$, making $0.01$ a clear dividing point. These thresholds effectively balance precision and applicability based on the observed data distributions.
@@ V-E Main results
The results of all methods are summarized in Table I. Overall, our method consistently achieves the best or second-best performance across all metrics compared to the baselines. Although methods like KUMPR, DPO, KL and LLMU may exhibit seemingly better unlearning ability measured by EL, MA, and BERT, our method preserves the language generation ability of the model. Compared to Opt, our model unlearns the original model better. Overall, our method achieves the best normalized scores across all model sizes, showing the best trade-off between unlearning and performance. Additionally, we observe several interesting phenomena:
Figure 3: Ablation results on hyperparameters ($\alpha,\beta$) for GPT-Neo 125M.
Figure 4: Case study for the comparison among various methods.
@@
First, the impact of unlearning becomes more pronounced with larger models, indicating that larger models have a higher tendency to memorize sensitive information, which our method effectively mitigates. Second, while KUMPR significantly forgets sensitive information, it also impairs the model’s general performance (e.g., PPL). In contrast, our ICU approach preserves the model’s core linguistic capabilities while erasing sensitive information, underscoring its advantage.
V-F Ablation study
@@
As discussed in Section IV-B, pair learning loss and KL-divergence loss are employed to ensure the model’s stable generative capability. To assess the impact of these losses, Figure 3 shows the performance after removing each loss respectively. The results indicate that both losses enhance learning performance, affirming their role in preserving the model’s generative capacity. Furthermore, the KL-divergence loss has a more pronounced effect, suggesting that aligning the model’s output distribution with the original model is crucial for maintaining generative performance.
V-G Human evaluation
We have also conducted human evaluation about the results of different methods. Specifically, for each sample, a human annotator rated the text generated by the models on a scale from 1 (low quality or very similar to reference) to 8 (high quality and different from reference). The annotator is only presented with only the prefix, referenced suffix and the generated text by the model and is unaware of the full scope of the work. The instruction and more details about human evaluation can be found in our available codes.
@@ V-G Human evaluation
The comparison between GPT and Human Evaluation scores can be found in Table II. From the results, human annotation and the GPT evaluation share the same trend, demonstrating the reliability of GPT-4 grading. For statistical test, we normalize the human score $s$ by $10\cdot\frac{s-1}{8-1}$, and the Pearson correlation is $0.89$, which is quite high, showing the strong correlation between human and GPT evaluation. Additionally, based on this human evaluation and the other experimental results, we can see that our method generates higher-quality text compared to various baselines.
TABLE II: Human annotation results on different methods.(Avg. Human scores / Avg. GPT scores.)
@@ V-H Case study
To provide a clearer comparison of our methods, we present a case study demonstrating the balance between learning and unlearning. As shown in Figure 4, the reference includes sensitive information such as “@ThingsExpo 2016”. Before unlearning, the original models (e.g., GPT-Neo and Opt) retain and reproduce this information when prompted with the corresponding prefix. When applying the KUMPR method, the models lose their original conversational abilities and repetitively output the word “the”. DPO method also produces sentences not coherent. KL and LLMU methods both repeat same sentence. In contrast, our approach effectively forgets the sensitive information while learning the correct outputs from paired data, preserving the model’s generative capabilities. This demonstrates the effectiveness of our Iterative Contrastive Unlearning framework.
Figure 5: Unlearning process illustration for the case provided in Figure 4.
TABLE III: ICU with different $\alpha$ and $\beta$ on GPT-Neo 125M. Our final parameter selection ($\alpha=0.5$ and $\beta=1.0$) is bolded.
$\alpha$ $\beta$ EL10 MA BERT Entropy Cls Avg. Dia Avg. Pile Wikitext
0.1 0.5 3.6 54.7 47.0 3.009 39.5 5.6 95.2 212.7
0.1 1.0 4.2 56.4 52.5 3.758 42.9 10.0 23.3 44.8
0.5 0.1 3.3 56.5 49.9 3.423 40.0 8.7 35.4 85.5
0.5 0.5 4.2 57.5 52.6 3.755 42.4 9.7 24.5 48.8
0.5 1.0 4.4 55.6 53.3 3.833 43.3 10.3 21.6 40.1
1.0 0.5 4.5 55.1 52.1 3.700 43.1 10.1 21.6 40.1
1.0 1.0 4.6 53.5 52.7 3.846 43.2 10.0 21.6 40.0
1.0 2.0 4.5 47.7 50.6 3.686 43.1 10.0 21.6 39.7
2.0 1.0 4.6 47.2 50.6 3.809 42.7 9.8 22.1 40.5
2.0 2.0 4.5 45.8 51.2 3.843 42.9 10.0 22.1 40.1
V-I Unlearning process analysis